# CoinFT: A Coin-Sized, Capacitive 6-Axis Force Torque Sensor for Robotic Applications

### v1

@@ III-D1 CoinFT Calibration
The 8 shear mode electrodes respond to inputs causing change in distance, $Fz$, $Mx$, and $My$, between the electrodes and the ground electrode (Fig. 4 (a)). While it is possible to calibrate the sensor using only the shear mode elecrodes, the redundancy added by the high signal-to-noise ratio (SNR) by the normal mode signals (Fig. 4 (a)) significantly increase calibration accuracy, as shown in Table I. As expected, the increase in accuracy is more significant for $Fz$, $Mx$, and $My$.
@@ III-D2 Sensitivity and Dynamic Range Tuning
Different robotic applications require varying sensor dynamic range, sensitivity, and accuracy. For instance, fine control of a fingertip-mounted haptic device [40] requires high sensitivity to small forces within a dynamic range under 4 N in the normal direction, whereas measuring the ground reaction force of quadruped robot feet may require a range of up to 350 N [41], with less emphasis on sensitivity to small forces. The design of the pillar layer can be adjusted to alter its mechanical properties, meeting the specific requirements of different applications.
@@ III-D3 Dynamic Response
A frequency response analysis has been conducted to investigate the mechanical bandwidth of CoinFT (Fig. 6). A dual-mode muscle lever (Aurora Scientific Inc., 309C) has been used to provide a sinusoidal force sweep either in the normal or shear directions from 0 Hz $\sim$ 120Hz. The force input ranged from 0 N $\sim$ 2 N in the normal direction and 0.1 N $\sim$ 1.1 N in the shear direction. As shown in Fig. 6 (a), the magnitude response of CoinFT in the normal direction is kept above 70.7% (-3dB) of the input signal across the entire swept frequency. The phase lag is negligible up to 2.5 Hz, and gradually increases with higher frequency input where, at 120 Hz, the phase lag is -34.6 °. The magnitude response in the shear direction (Fig. 6 (b)) starts to deteriorate at 2.45Ḣz and reaches 70.7% (-3dB) of the input signal at 97 Hz. At that point, the phase delay is -94.2°. The cause of the phase lag is likely due to the innate hysteresis of silicone rubber material, as shown in Fig. 4 (d). These results show that the mechanical bandwidth of CoinFT is approximately 97 Hz.
@@ IV-A Drone with CoinFT
Our drone is a 127 mm frame quadrotor with an UpBoard companion computer and a PixRacer flight controller (Fig. 8). CoinFT is attached at the distal tip of a custom mount, which is connected in series with the proximal mount through a telescoping mechanism with rubber bands (Fig. 8(c)). The compliant element in series help reduce impact upon contact. CoinFT is connected to the UpBoard computer through a USB port. A socket is attached at the top of CoinFT which allows mounting different end-effectors for different tasks, such as a round tip for general contact or a package of electronics (Fig. 8(a) and (b)). The electronics package consists of an ESP32 microcontroller that uses four 12 V batteries in parallel to control a strip of LED to signal successful package deployment. Pressure sensitive adhesives are attached at the corners on the top side of the package. Since the package is not mechanically attached to the socket, the package will attach to a surface once sufficient pressure provides the necessary adhesion.
@@ IV-B Attitude Controller for Contact
We assume the quadrotor has access to reliable full state information. While technologies exist to achieve this with on-board only sensors, implementing them falls outside the scope of this paper, so we utilize motion capture to assist state estimation by providing the drone with position information. We also provide the quadrotor with a minimum snap trajectory spline [43] that passes through the contact surface. This allows us to take a range of heights within which we are confident that the contact surface exists and transform it into a sequence of desired positions ($\boldsymbol{p_{des}}$), velocities ($\boldsymbol{v_{des}}$) and orientations ($\boldsymbol{q_{des}}$), for our controller to track while it searches for contact.
@@
To account for the drone flight envelope including both in-contact and out-of-contact flight we use a switching, cascaded PID controller to advance from currently observed states $\begin{bmatrix}\boldsymbol{p_{obs}}&\boldsymbol{v_{obs}}&\boldsymbol{q_{obs}}\end{bmatrix}$ to desired attitude commands $\boldsymbol{u_{cmd}}$. We start by taking the position and velocity error:
@@
where $F_{des}$ is the desired force, $m$ is the mass of the drone, $g$ is the gravity vector, and ($\boldsymbol{K_{p}},\boldsymbol{K_{v}}$) are positive definite gain matrices that switch between in-contact gains and out-of-contact gains:
@@
Using the quaternion to rotation matrix to basis vector mapping $\boldsymbol{q_{i}}\Longleftrightarrow\boldsymbol{R_{i}}\Longleftrightarrow\begin{bmatrix}\boldsymbol{x_{i}}&\boldsymbol{y_{i}}&\boldsymbol{z_{i}}\end{bmatrix}$ on both the desired orientation ($\boldsymbol{q_{des}}$) and commanded orientation ($\boldsymbol{q_{cmd}}$), we can compute the latter by doing:
@@
where $k_{f}$ is the thrust normalization coefficient. This is then fed into a state machine (Algorithm 1) for thrust control that switches modes during contact events and tracks desired contact forces using CoinFT feedback. The two key features of the state machine is a linear search function for the contact force and a PID tracking controller. The former enables graceful transitions when the initial desired contact force is large while the latter not only enables us to hold but also track a changing desired contact force. Combined, these allow for more consistent and safer performance. Given an observed force on CoinFT ($f_{oc}$), desired force on CoinFT ($f_{dc}$) and current time ($t_{cr}$) we perform the following to get the commanded normalized thrust. Assuming compressive force, we use the absolute values of the CoinFT reading in the normal direction (Algorithm 1).
@@
The CoinFT feedback data lets us use the state machine to dynamically switch the gain matrices $\boldsymbol{K_{p}},\boldsymbol{K_{v}}$ where we use the out of contact gains when in FREE mode and in contact gains otherwise.
@@ IV-C Force Controlled Deployment of Objects on Environment Surfaces
To test the designed force controller and perform electronics package delivery tasks, we conducted flight tests. The test setup includes a laser-cut acrylic plate horizontally attached to a structure made of extruded aluminum frames (Fig. 10 (a)), simulating arbitrary environmental surfaces. The integrated system consists of a ground control station (PC), motion capture system (OptiTrack, NaturalPoint, Inc.), and a drone equipped with a companion computer, flight controller, IMU, and a CoinFT sensor (Fig. 9). The proposed attitude controller runs on the companion computer, providing control inputs to the flight controller, which then converts these inputs into PWM signals for the four motors.
@@
Using the force controller, the drone can attach electronics packages to environmental surfaces such as a tree branch by applying a precise range of force—sufficient to complete the task but not so excessive as to damage the environment, or expend more energy than required. The electronics package, equipped with pressure-sensitive adhesives and weighing 95 g (Fig. 8 (b)), is mounted on the drone. Adequate pressure must be applied for the package to adhere securely to a surface. The drone initially hovers below the horizontal surface and ascends until contact is detected. Upon contact, the contact force ramps up to a commanded value of 0.7 N, which registers as approximately 1.7 N in CoinFT due to the package’s weight (Fig. 11 (a)). The drone breaks contact and descends while still sensing approximately 1 N in CoinFT, detecting that the package has not adhered properly (Fig. 11 (b)). The drone then autonomously increases the desired contact force to 5 N (Fig. 11 (c)). After applying the higher force, the drone no longer senses the package’s weight through CoinFT and sends an ON command to the electronics package via WiFi, activating the LEDs to signal task success (Fig. 11 (d)). These precisely controlled interactions are valuable for attaching objects, such as sensor packages, to potentially fragile surfaces like tree branches or for performing general force-controlled tasks, such as cleaning windows on tall buildings.
@@ V Discussion & Future Work
While this paper focuses primarily on demonstrating CoinFT’s capabilities on drone platforms, the sensor’s versatility extends to a wide range of other robotic applications. CoinFT has already been used in wearable haptic devices, where it measures and controls the contact force between the device and the human subject. For instance, Sarac et al. [44] used CoinFT to reveal that people perceive different intensities when the same force is applied in normal versus shear directions. Yoshida et al. [7] employed CoinFT to ensure consistency in user studies by initializing the haptic device tactor based on contact force rather than position. This approach is crucial for accommodating variations in users’ body shapes, as position-based initialization does not consistently deliver uniform haptic feedback. CoinFT’s slim and compact design also makes it highly suitable for wearable fingertip devices [40].
CoinFT has also been directly attached under the fingertips of a commercial anthropomorphic robot hand [45] for learning force-informed actions from kinesthetic demonstrations. In such setup, it can sense contact force for controlled, forceful interaction or track the contact point by utilizing force and torque information for intrinsic tactile sensing [46]. When connected via I2C communication, multiple CoinFTs can be integrated to sensorize all fingers, enhancing the robot hand’s ability to perform fine-manipulation tasks. Its low cost and robustness against incidental impacts make CoinFT particularly desirable for robot hands operating in unstructured and cluttered environments.
@@
While CoinFT is versatile and suitable for a wide range of robotic applications, it is not intended as a direct replacement for existing commercial sensors like the Gamma (ATI Industrial Automation). The mechanical bandwidth of CoinFT is an order of magnitude lower than that of the Gamma (1 kHz), making it less capable of detecting highly dynamic contacts. Additionally, the Gamma offers a finer force resolution even with its high dynamic response capabilities. Although CoinFT is highly robust against large compressive impulses, it is less resilient to large tensile impacts, as the bonding between the silicone rubber pillars and the PCB may fail under such conditions.
Other challenges must be considered when using CoinFT. While shielding has been implemented to prevent the coupling of stray capacitance, the surface-mounted electrodes of the chip remain partially exposed, which can result in imperfect shielding. Furthermore, the silicone rubber pillars may creep under long-term sustained loads, altering the mechanical characteristics of the pillar array and necessitating re-calibration. The current thin PCB might not be entirely rigid and could deform out of plane under significant local pressure, potentially reducing calibration accuracy. This issue, however, can be mitigated by using a slightly thicker PCB or bonding a thin layer of stiff material. Additionally, variations in ambient temperature may cause slight deformations in the sensing electrodes or the internal electronics of the chip, leading to drift in the capacitance signal. Despite these limitations, CoinFT remains valuable for many robotic applications where such high precision and bandwidth in a 6-axis F / T sensor is not essential, and where robustness and low cost are prioritized.
@@ VI Conclusion
We demonstrated the utility of CoinFT by integrating it with a drone to perform force-controlled tasks, such as attaching an electronics package to environmental surfaces. This capability is an example of CoinFT’s potential to democratize the use of 6-axis F / T sensors across a broad range of robotic platforms, enabling more widespread adoption in applications that demand robustness and cost-effectiveness without sacrificing essential functionality.
@@ ACKNOWLEDGEMENTS
This work was supported in part by Meta Reality Labs Research and the Kwanjeong Fellowship. The authors thank Claire Chen, Kyle Yoshida, Crystal Winston, Mine Sarac, and Alice Wu for exciting and insightful discussions about CoinFT applications.

### v2

@@ III-D1 CoinFT Calibration
The 8 shear mode electrodes respond to inputs causing change in distance, $Fz$, $Mx$, and $My$, between the electrodes and the ground electrode (Fig. 4 (a)). While it is possible to calibrate the sensor using only the shear mode elecrodes, the redundancy added by the high signal-to-noise ratio (SNR) by the normal mode signals (Fig. 4 (a)) significantly increases calibration accuracy, as shown in Table I. As expected, the increase in accuracy is more significant for $Fz$, $Mx$, and $My$.
@@ III-D2 Sensitivity and Dynamic Range Tuning
Different robotic applications require varying sensor dynamic range, sensitivity, and accuracy. For instance, fine control of a fingertip-mounted haptic device [43] requires high sensitivity to small forces within a dynamic range under 4 N in the normal direction, whereas an instrumented probe for ultrasound scans might require sustained loads of 20 $\sim$ 40 N [44] and an instrumented tool for physiotherapy manipulation might require 100 $\sim$ 200 N—with correspondingly less emphasis on sensitivity to small forces [45]. The design of the pillar layer can be adjusted to alter its mechanical properties, meeting the specific requirements of different applications.
@@ III-D3 Dynamic Response
A frequency response analysis has been conducted to investigate the mechanical bandwidth of CoinFT (Fig. 6). A dual-mode muscle lever (Aurora Scientific Inc., 309C) has been used to provide a sinusoidal force sweep either in the normal or shear directions from 0 Hz $\sim$ 120Hz. The force input ranged from 0 N $\sim$ 2 N in the normal direction and 0.1 N $\sim$ 1.1 N in the shear direction. As shown in Fig. 6 (a), the magnitude response of CoinFT in the normal direction is kept above 70.7% (-3dB) of the input signal across the entire swept frequency. The phase lag is negligible up to 2.5 Hz, and gradually increases with higher frequency input where, at 120 Hz, the phase lag is -34.6 °. The magnitude response in the shear direction (Fig. 6 (b)) starts to deteriorate at 2.45Ḣz and reaches 70.7% (-3dB) of the input signal at 97 Hz. At that point, the phase delay is -94.2°. The cause of the phase lag is likely due to the innate hysteresis of silicone rubber material, visible in Fig. 4 (d). These results show that the mechanical bandwidth of CoinFT is approximately 97 Hz.
@@ IV-A Drone with CoinFT
Our drone is a 127 mm frame quadrotor with an UpBoard companion computer and a PixRacer flight controller (Fig. 8). CoinFT is attached at the distal tip of a custom mount, which is connected in series with the proximal mount through a telescoping mechanism with rubber bands (Fig. 8(c)). The compliant element in series helps to reduce impact upon contact and simplify control. CoinFT is connected to the UpBoard computer through a USB port. A socket is attached at the top of CoinFT which allows mounting different end-effectors for different tasks, such as a round tip for general contact or a package of electronics (Fig. 8(a) and (b)). The electronics package consists of an ESP32 microcontroller that uses four 12 V batteries in parallel to control a strip of LEDs to signal successful package deployment. Pressure sensitive adhesives are attached at the corners on the top side of the package. Since the package is not mechanically attached to the socket, the package will attach to a surface once sufficient pressure provides the necessary adhesion.
@@ IV-B Attitude Controller for Contact
Given that the focus here is control with respect to contact, we simplify the navigation aspect of the control problem by providing the drone with accurate estimates of its pose and orientation via an external motion capture system. We then use this as feedback to track a minimum snap trajectory spline [47] to a region near the contact surface. Crucially though, we do not provide the drone with the actual height of the contact surface. Instead, our policy will search through a range of heights within which we are confident that the contact surface exists and transform it into a sequence of desired positions ($\boldsymbol{p_{des}}$), velocities ($\boldsymbol{v_{des}}$) and orientations ($\boldsymbol{q_{des}}$), for our controller to track while it searches for contact.
@@
To account for the drone flight envelope including both in-contact and out-of-contact flight, we use a switching, cascaded PID controller to advance from currently observed states $\begin{bmatrix}\boldsymbol{p_{obs}}&\boldsymbol{v_{obs}}&\boldsymbol{q_{obs}}\end{bmatrix}$ to desired attitude commands $\boldsymbol{u_{cmd}}$. We start by taking the position and velocity error:
@@
where $\boldsymbol{F_{des}}$ is the desired force, $m$ is the mass of the drone, $\boldsymbol{g}$ is the gravity vector, and ($\boldsymbol{K_{p}},\boldsymbol{K_{v}}$) are positive definite gain matrices that switch between in-contact gains and out-of-contact gains:
@@
Using the quaternion to rotation matrix to basis vector mapping $\boldsymbol{q_{i}}\Longleftrightarrow\boldsymbol{R_{i}}\Longleftrightarrow\begin{bmatrix}\boldsymbol{x_{i}}&\boldsymbol{y_{i}}&\boldsymbol{z_{i}}\end{bmatrix}$ on both the desired orientation ($\boldsymbol{q_{des}}$) and commanded orientation ($\boldsymbol{q_{cmd}}$), we can compute the latter:
@@
where $k_{f}$ is the thrust normalization coefficient that can be empirically approximated from prior flights. This is then fed into a state machine (Algorithm 1) for thrust control that switches modes during contact events and tracks desired contact forces using CoinFT feedback. The two key features of the state machine is a linear search function for the contact force and a PID tracking controller. The former enables graceful transitions when the initial desired contact force is large while the latter not only enables us to hold but also track a changing desired contact force. Combined, these allow for more consistent and safer performance. Given an observed force on CoinFT ($f_{oc}$), desired force on CoinFT ($f_{dc}$) and current time ($t_{cr}$) we perform the steps in Algorithm 1 to get the commanded normalized thrust. Assuming compressive force, we use the absolute values of the CoinFT reading in the normal direction.
@@
The CoinFT feedback data lets us use the state machine to dynamically switch the gain matrices $\boldsymbol{K_{p}},\boldsymbol{K_{v}}$ where we use the out of contact gains when in FREE state and in contact gains otherwise (Algorithm 1).
@@ IV-C Force Controlled Deployment of Objects on Environment Surfaces
To test the designed force controller and perform electronics package delivery tasks, we conducted flight tests. The test setup includes a laser-cut acrylic plate horizontally attached to a structure made of extruded aluminum frames (Fig. 10 (a)), simulating an environmental surface. The integrated system consists of a ground control station (PC), motion capture system (OptiTrack, NaturalPoint, Inc.), and a drone equipped with a companion computer, flight controller, IMU, and a CoinFT sensor (Fig. 9). The proposed attitude controller runs on the companion computer, providing control inputs to the flight controller, which then converts these inputs into PWM signals for the four motors.
@@
Using the force controller, the drone can attach electronics packages to environmental surfaces such as a tree branch by applying a range of force—sufficient to complete the task but not sufficient to damage the environment, or expend more energy than required. The electronics package, equipped with pressure-sensitive adhesives and weighing 95 g (Fig. 8 (b)), is mounted on the drone. Adequate pressure must be applied for the package to adhere securely to a surface. The drone initially hovers below the horizontal surface and ascends until contact is detected. Upon contact, the contact force ramps up to a commanded value of 0.7 N, which registers as approximately 1.7 N in CoinFT due to the package’s weight (Fig. 11 (a)). The drone breaks contact and descends while still sensing approximately 1 N in CoinFT, detecting that the package has not adhered properly (Fig. 11 (b)). The drone then autonomously increases the desired contact force to 5 N (Fig. 11 (c)). After applying the higher force, the drone no longer senses the package’s weight through CoinFT and sends an ON command to the electronics package via WiFi, activating the LEDs to signal task success (Fig. 11 (d)). These precisely controlled interactions are valuable for attaching objects, such as sensor packages, to potentially fragile surfaces like tree branches or for performing general force-controlled tasks, such as cleaning windows on tall buildings.
@@ V Discussion & Future Work
While this paper focuses primarily on demonstrating CoinFT’s capabilities on drone platforms, the sensor’s versatility extends to a wide range of other robotic applications. CoinFT has already been used in wearable haptic devices, where it measures and controls the contact force between the device and the human subject. Sarac et al. [48] used CoinFT to reveal that people perceive different intensities when the same force is applied in normal versus shear directions. Yoshida et al. [7] employed CoinFT to ensure consistency in user studies by initializing the haptic device tactor based on contact force rather than position. This approach is crucial for accommodating variations in users’ body shapes, as position-based initialization does not consistently deliver uniform haptic feedback. CoinFT’s slim and compact design also makes it highly suitable for wearable fingertip devices [43].
CoinFT has also been directly attached to the fingertips of a commercial anthropomorphic robot hand [49] for learning force-informed actions from kinesthetic demonstrations. In such setup, it can sense contact force for controlled, forceful interaction or track the contact point by utilizing force and torque information for intrinsic tactile sensing [50]. When connected via I2C communication, multiple CoinFTs can be integrated to sensorize all fingers at the same 360 Hz sampling rate, enhancing the robot hand’s ability to perform fine-manipulation tasks. Its low cost and robustness against incidental impacts make CoinFT particularly desirable for robot hands operating in unstructured and cluttered environments.
@@
While CoinFT is versatile and suitable for a wide range of robotic applications, it is not intended as a direct replacement for existing commercial sensors like the Gamma (ATI Industrial Automation). The mechanical bandwidth of CoinFT is an order of magnitude lower than that of the Gamma (1 kHz), making it less capable of detecting highly dynamic contacts. Additionally, the Gamma offers a finer force resolution even with its high dynamic response capabilities. Although CoinFT is highly robust against large compressive impulses, it is less resilient to large tensile loads, as the bonding between the silicone rubber pillars and the PCB may fail under such conditions.
Other challenges must be considered when using CoinFT. While shielding has been implemented to prevent the coupling of stray capacitance, the surface-mounted electrodes of the chip remain partially exposed, which can result in imperfect shielding. Furthermore, the silicone rubber pillars may creep under long-term sustained loads, altering the mechanical characteristics of the pillar array and necessitating re-calibration. For applications that require it, one could assemble CoinFT with a low-creep silicone elastomer [52] of similar stiffness.
The current thin PCB might not be entirely rigid and could deform out of plane under significant local pressure, potentially reducing calibration accuracy. This issue, however, can be mitigated by using a slightly thicker PCB or bonding a thin layer of stiff material. Additionally, variations in ambient temperature may cause slight deformations in the sensing electrodes or the internal electronics of the chip, leading to drift in the capacitance signal. Despite these limitations, CoinFT remains valuable for many robotic applications where such high precision and bandwidth in a 6-axis F / T sensor is not essential, and where robustness and low cost are prioritized.
@@ VI Conclusion
We demonstrated the utility of CoinFT by integrating it with a drone to perform force-controlled tasks, such as attaching an electronics package to environmental surfaces. This capability is an example of CoinFT’s potential to democratize the use of 6-axis F / T sensors across a broad range of robotic platforms, enabling more widespread adoption in applications that demand robustness and cost-effectiveness without sacrificing essential functionality. To promote this use of CoinFT technology, we have published solid models, fabrication instructions and software as an open-source design at https://hojung-choi.github.io/coinft.github.io.
@@ ACKNOWLEDGEMENTS
Meta Reality Labs Research, Toyota Research Institute, and the Kwanjeong Fellowship provided funds to support this work. The authors thank Claire Chen, Kyle Yoshida, Crystal Winston, Mine Sarac, and Alice Wu for exciting and insightful discussions about CoinFT applications.